# UniBERTs: Adversarial Training for Language-Universal Representations

### v1

@@ Abstract
This paper presents UniBERT, a compact multilingual language model that leverages an innovative training framework integrating three components: masked language modeling, adversarial training, and knowledge distillation. Pre-trained on a meticulously curated Wikipedia corpus spanning 107 languages, UniBERT is designed to reduce the computational demands of large-scale models while maintaining competitive performance across various natural language processing tasks. Comprehensive evaluations on four tasks—named entity recognition, natural language inference, question answering, and semantic textual similarity—demonstrate that our multilingual training strategy enhanced by an adversarial objective significantly improves cross-lingual generalization. Specifically, UniBERT models show an average relative improvement of 7.72% over traditional baselines, which achieved an average relative improvement of only 1.17%, with statistical analysis confirming the significance of these gains (p-value = 0.0181). This work highlights the benefits of combining adversarial training and knowledge distillation to build scalable and robust language models, thereby advancing the field of multilingual and cross-lingual natural language processing.
@@ 5.5 Discussion
The results indicate that the multilingual training strategy, in conjunction with the adversarial objective, has a pronounced impact on overall performance. Figure 3 illustrates the relative improvements achieved by each model when transitioning from individual language fine-tuning (denoted as "each") to joint multilingual fine-tuning (denoted as "all") [^6] Since the score on the STS lied between 0 and 1, we normalized it before calculating the mean relative improvement.. mBERT, XLM-RoBERTa, and Distil-mBERT versions exhibit a moderate average relative improvements of approximately 1.17%: 1.51% for Distil-mBERT-base, 0.62% for mBERT-base, 2.11% for XLM-RoBERTa-baase and 0.44% for XLM-RoBERTa-large. In contrast, UniBERT models indicate significantly greater gains, with an average relative improvement of 7.72%: 3.76% for UniBERT-small, 10.33% for uniBERT-XSmall and 9.08% for UniBERT-XXSmall.
This significant difference suggests that the adversarial training component, which promotes the learning of language-agnostic representations, is particularly effective when combined with joint multilingual training. For instance, the UniBERT-XSmall model improves from 40.20 in the "each" configuration to 44.35 in the "all" configuration, and the UniBERT-XXSmall model similarly increases from 34.51 to 37.64. Such improvements are notably higher than those observed in larger models, where the increase is relatively modest.

### v2

# UniBERT: Adversarial Training for Language-Universal Representations

@@ Abstract
This paper presents UniBERT, a compact multilingual language model that uses an innovative training framework that integrates three components: masked language modeling, adversarial training, and knowledge distillation. Pre-trained on a meticulously curated Wikipedia corpus spanning 107 languages, UniBERT is designed to reduce the computational demands of large-scale models while maintaining competitive performance across various natural language processing tasks. Comprehensive evaluations on four tasks—named entity recognition, natural language inference, question answering, and semantic textual similarity—demonstrate that our multilingual training strategy enhanced by an adversarial objective significantly improves cross-lingual generalization. Specifically, UniBERT models show an average relative improvement of 7.72% over traditional baselines, which achieved an average relative improvement of only 1.17%, and statistical analysis confirms the significance of these gains (p-value = 0.0181). This work highlights the benefits of combining adversarial training and knowledge distillation to build scalable and robust language models, thus advancing the field of multilingual and cross-lingual natural language processing.
@@ 5.5 Discussion
The results indicate that the multilingual training strategy, in conjunction with the adversarial objective, has a pronounced impact on overall performance. Figure 3 illustrates the relative improvements achieved by each model when transitioning from individual language fine-tuning (denoted as "each") to joint multilingual fine-tuning (denoted as "all") [^6] Since the score on the STS is between 0 and 1, we normalized it before calculating the mean relative improvement.. mBERT-base, Distil-mBERT-base, and the XLM-RoBERTa versions exhibit a moderate average relative improvement of approximately 1.17%: 1.51% for Distil-mBERT-base, 0.62% for mBERT-base, 2.11% for XLM-RoBERTa-base, and 0.44% for XLM-RoBERTa-large. In contrast, UniBERT models indicate significantly greater gains, with an average relative improvement of 7.72%: 3.76% for UniBERT-small, 10.33% for uniBERT-XSmall, and 9.08% for UniBERT-XXSmall.
This significant difference suggests that the adversarial training component, which promotes the learning of language-agnostic representations, is particularly effective when combined with joint multilingual training. For example, the UniBERT-XSmall model improves from 40.20% in the "each" configuration to 44.35% in the "all" configuration, and the UniBERT-XXSmall model similarly increases from 34.51% to 37.64%. Such improvements are notably higher than those observed in larger models, where the increase is relatively modest (e.g., XLM-RoBERTa-base and XLM-RoBERTa-large improve by less than 1.2 and 0.3 percentage points, respectively).
@@ 7 Limitations
Acknowledgements
This work was supported by the National University of Science and Technology POLITEHNICA Bucharest through the PubArt program. DCC is funded by the National Program for Research of the National Association of Technical Universities (GNAC ARUT 2023).